OpenStudy (anonymous):

4log^x – 6log^(x + 2) (Assuming the bases are 10)

OpenStudy (anonymous):

whats this for ^

OpenStudy (anonymous):

Alg 2

OpenStudy (anonymous):

I'm just supposed to combine them into one log.

OpenStudy (anonymous):

okay first\[a \log b=\log b ^a\]

OpenStudy (anonymous):

so in our case take 4 and 6 to be the power of inside expressions

OpenStudy (anonymous):

So would it look like log^x^4 - log^(x + 2)^6?

OpenStudy (anonymous):

yes now we use \[\log a-\log b=\log \frac{a}{b}\]

OpenStudy (anonymous):

log (x^4)/(x+ 2)^6 right?

OpenStudy (anonymous):

yes\[\log \frac{x^4}{(x+2)^6}\] correct

OpenStudy (anonymous):

Is that the final form or can it be simplified from there

OpenStudy (anonymous):

final

OpenStudy (anonymous):

Thanks!